[The following information applies to the questions displayed below.]

Super Saver Groceries purchased store equipment for $31,000. Super Saver estimates that at the end of its 10-year service life, the equipment will be worth $4,000. During the 10-year period, the company expects to use the equipment for a total of 10,000 hours. Super Saver used the equipment for 1,500 hours the first year.

Required:

Calculate depreciation expense of the equipment for the first year, using each of the following methods. (Do not round your intermediate calculations.)

1. Straight-line. Depreciation expense

2. Double-declining-balance. Depreciation expense

3. Activity-based. Depreciation expense
Answer #1

Straight line = (Cost - Salvage value) /useful life

= (31,000-4000)/10

= 2700

Double declining rate = 200/useful life = 200/10 = 20%

= 31,000*10%

= 3100

Activity based

= (31,000-4,000)/10,000 * 1500

= 4050